OpenStudy (anonymous):

I need some help with 7 Adding and Subtracting Radical Expressions problems. :)

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form.\[\sqrt{16y^3}-\sqrt{y^3}\]

OpenStudy (anonymous):

Please explain how you got the answer, and I'll try some of te others on my own- I may need some help though :)

OpenStudy (anonymous):

So would the answer be \[8\sqrt{8}\]?

OpenStudy (anonymous):

oops, I meant \[4\sqrt{2}\]

OpenStudy (anonymous):

Oh, okay :)

OpenStudy (anonymous):

So then it beomes \[4y^2\sqrt{y}\]?

OpenStudy (anonymous):

\[3y \sqrt{y}\] then... Wow I'm so sorry, my teacher wouldn't explain it well.

OpenStudy (anonymous):

I haven't gotten to multiplying and dividing yet, but that's next :)

OpenStudy (anonymous):

Oh, I see! I'm going to try the next problem, but I might need some help... It's Find the exact value of the radical expression in simplest form. \[5\sqrt{2}+\sqrt{2}\]